If the market for bread is described by the following equations Qd=210-0,6P and Qs=-96+1,1P what is the equilibrium price for bread?  

1
Expert's answer
2021-06-11T11:47:37-0400
"210-0.6p=-96+1.1p"

"1.1p+0.6p=210+96"

"1.7p=306"

"p=180"